# Theory of Mind abilities of Large Language Models in Human-Robot Interaction : An Illusion?

### 10 Jan 2024 v1

@@ C.4. USAR Domain
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There are two med-kits: a) medkit1 - heavier & lies closer to the room where CommX is, and b) medkit2 - lighter & lies across the hallway close to the room where a patient is located. The observer (you) has the top-view of this setting, and do not know about the properties of the med-kits. Goal : Agent has to pickup a med-kit and hand it over to CommX in the shortest plan possible. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot picks up medkit2 and hands it over to CommX. Question 1 : Imagine you are the human observer in this case. Would you find such a plan explicable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the agent should have picked up medkit1 and handed it over to CommX. 2) Because the agent should pick up the medkit which is farther away. 3) Because no plan is explicable.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks up the medkit and turns LEFT. Question 1 : Would you find such a partial plan legible? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the agent is going to the patient room directly. 2) Because the agent is going to CommX to handover the medkit. 3) Because CommX can not take medkit to the patient room.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway, and there are two paths to a patient room from there (one on the LEFT, and one on the RIGHT). The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and take it to the patient room. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up the medkit, and takes one step LEFT. Question 1 : Would you find such a partial plan predictable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot will go to the patient room from the LEFT path. 2) Because the robot will go to the patient room from the RIGHT path. 3) Because all paths are predictable.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the medkit and is now at an equal distance from the patient room, and CommX. Question 1 : Would you find such a partial plan obfuscatory? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot can possibly go to either the patient room, or to CommX. 2) Because the robot will only go to the patient room. 3) Because the robot will only go to CommX.
@@ C.5. Package Delivery Domain
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port that can either be acquired on the port, or else loaded on the ship by the robot. Goal: A crate needs to be loaded on the ship as fast as possible. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot first acquires a crate on the port, and then loads a crate on the ship. Question 1 : Imagine you are the human observer in this case. Would you find such a plan explicable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot should have loaded the crate first to achieve the goal, and then acquired another crate on the port. 2) Because the human observer can find out about the identity of the loaded crate on the ship. 3) Because no plan is explicable.
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port. Goal: The robot can either pick and acquire a package on the port, or else pick and load it on the ship. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks up the package and brings it closer to the ship. Question 1 : Would you find such a partial plan legible? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot is first going to load the package on the ship. 2) Because the robot is going to acquire the package on the port. 3) Because the human observer knows about the identity of the package.
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, the human observer only gets to know the identity of the crate when it is either acquired at the port or loaded on the ship. Initial state: There are packages at the port. Goal: The robot has to reveal the identity of a package to the human. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up a package and goes towards the port. Question 1 : Would you find such a partial plan predictable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot will now acquire the package at the port to achieve the goal. 2) Because the robot will now load the package on the ship to achieve the goal. 3) Because the robot will directly show the package to the human.
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port. Goal: The robot can either pick and acquire a package on the port, or else pick and load it on the ship. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the package and holds it between the port and the ship. Question 1 : Would you find such a partial plan obfuscatory? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot can possibly either acquire the package on the port or load it on the ship. 2) Because the robot will only acquire the package on the port. 3) Because the robot will only load the package on the ship.
Appendix D LLM Prompts: Uninformative Context
@@ D.1. Fetch Robot Domain
Description: Fetch is a robot that can carry objects (pick / place) and move from one location to another. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Goal: The robot has to take the block b1 to location loc2. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot tucks its arms and lowers itself. It then picks up the block b1. Finally, the robot moves to location loc2. The robot has a label saying ’Explicable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a plan explicable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) The plan is not explicable because the robot tucked its arm before moving, which seems unnecessary. 2) The plan is not explicable because the robot picked up block b1 which seems unnecessary. 3) The plan is not explicable because no plan is explicable.
@@
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) The plan is legible because the robot made its goal of taking the block to location loc2 more legible. 2) The plan is legible because the robot made its goal of picking up the block more legible. 3) The plan is legible because all robot plans are legible.
@@
Description: Fetch is a robot that can carry objects from one location to another. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location locX can be reached from location loc1 using two paths, one by taking five steps to left and the other one by taking five steps to right. Goal: The robot has to pick the block b1 and take it to location locX. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step towards left. The robot has a label saying ’Not a Predictable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan predictable? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) because the robot will go to locX from left path. 2) because the robot will go to locX from right path. 3) because all paths are predictable.
@@
Description: Fetch is a robot that can carry objects from one location to another. Fetch robot’s design requires it to tuck its arms and lower its torso or crouch before moving. There are three locations: loc1, loc2, and loc3. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the block b1. At loc1 is the agent is at a distance of 10 steps from loc2 and loc3. It takes 3 steps forward and is 7 steps away from loc2 as well as loc3. The robot has a label saying ’Not an Obfuscatory Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan obfuscatory? Give your answer as Yes or No only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) because the robot can possibly go to either location loc2 or location loc3. 2) because the robot will only go to location loc2. 3) because all plans are obfuscatory.

### 17 Jan 2024 v2

@@ C.4. USAR Domain
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There are two med-kits: a) medkit1 - heavier & lies closer to the room where CommX is, and b) medkit2 - lighter & lies across the hallway close to the room where a patient is located. The observer (you) has the top-view of this setting, and do not know about the properties of the med-kits. Goal : Agent has to pickup a med-kit and hand it over to CommX in the shortest plan possible. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot picks up medkit2 and hands it over to CommX. The robot has a label saying ’Explicable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a plan explicable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the agent should have picked up medkit1 and handed it over to CommX. 2) Because the agent should pick up the medkit which is farther away. 3) Because no plan is explicable.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks up the medkit and turns LEFT. The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the agent is going to the patient room directly. 2) Because the agent is going to CommX to handover the medkit. 3) Because CommX can not take medkit to the patient room.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway, and there are two paths to a patient room from there (one on the LEFT, and one on the RIGHT). The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and take it to the patient room. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up the medkit, and takes one step LEFT. The robot has a label saying ’Not a Predictable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan predictable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot will go to the patient room from the LEFT path. 2) Because the robot will go to the patient room from the RIGHT path. 3) Because all paths are predictable.
@@
Description : In a typical Urban Search and Rescue (USAR) setting, there is a building with interconnected rooms and hallways. There is a human commander CommX, and a robot agent acting in the environment. Both the agents can move around and pickup/drop-off or handover med-kits to each other. CommX can only interact with med-kits light in weight, but the robot agent can interact with heavy med-kits too. Initial State : There is a medkit located in the middle of the hallway. A patient is located on the LEFT of medkit and CommX is located on the RIGHT of medkit. The observer (you) has the top-view of this setting. Goal : Agent has to pickup medkit and can either take it to the patient room OR handover to CommX. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the medkit and is now at an equal distance from the patient room, and CommX. The robot has a label saying ’Not an Obfuscatory Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan obfuscatory? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot can possibly go to either the patient room, or to CommX. 2) Because the robot will only go to the patient room. 3) Because the robot will only go to CommX.
@@ C.5. Package Delivery Domain
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port that can either be acquired on the port, or else loaded on the ship by the robot. Goal: A crate needs to be loaded on the ship as fast as possible. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot first acquires a crate on the port, and then loads a crate on the ship. The robot has a label saying ’Explicable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a plan explicable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot should have loaded the crate first to achieve the goal, and then acquired another crate on the port. 2) Because the human observer can find out about the identity of the loaded crate on the ship. 3) Because no plan is explicable.
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port. Goal: The robot can either pick and acquire a package on the port, or else pick and load it on the ship. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks up the package and brings it closer to the ship. The robot has a label saying ’Not a Legible Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot is first going to load the package on the ship. 2) Because the robot is going to acquire the package on the port. 3) Because the human observer knows about the identity of the package.
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, the human observer only gets to know the identity of the crate when it is either acquired at the port or loaded on the ship. Initial state: There are packages at the port. Goal: The robot has to reveal the identity of a package to the human. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks up a package and goes towards the port. The robot has a label saying ’Not a Predictable Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan predictable? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot will now acquire the package at the port to achieve the goal. 2) Because the robot will now load the package on the ship to achieve the goal. 3) Because the robot will directly show the package to the human.
@@
Description: Consider a situation where there is a robot which manages a shipping port, and a human observer (you) who is the supervisor that has sensors or subordinates at the port who provide partial information about the nature of activity being carried out at the port. For instance, when a specific crate is loaded onto the ship, the observer finds out about the identity of the loaded crate. The observer knows the initial inventory at the port, but when new cargo is acquired by the port, the observer’s sensors reveal only that more cargo was received; they do not specify the numbers or identities of the received crates. Initial state: There are packages at the port. Goal: The robot can either pick and acquire a package on the port, or else pick and load it on the ship. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the package and holds it between the port and the ship. The robot has a label saying ’Not an Obfuscatory Agent’. The human observer cannot read. The human observer looks at the label. The human observer has not seen the robot before. Question 1 : Imagine you are the human observer. Would you find such a partial plan obfuscatory? Give your answer as ‘Yes’ or ‘No’ only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the robot can possibly either acquire the package on the port or load it on the ship. 2) Because the robot will only acquire the package on the port. 3) Because the robot will only load the package on the ship.
Appendix D LLM Prompts: Inconsistent Belief
@@ D.1. Fetch Robot Domain
Description: Fetch is a robot that can carry objects (pick / place) and move from one location to another. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Goal: The robot has to take the block b1 to location loc2. The human observer can not see the robot acting in the environment. Definition : Plan Explicability means whether the plan / robot behavior is an expected behavior according to the human observer (you). If you look at the robot behavior and find that some actions are unnecessary or not required, then the behavior is inexplicable. Plan : The robot tucks its arms and lowers itself. It then picks up the block b1. Finally, the robot moves to location loc2. Question 1 : Imagine you are the human observer. Would you find such a plan explicable? Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is not explicable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.
@@
Description: Fetch is a robot that can carry objects from one location to another. There are three locations: loc1, loc2, and loc3 where the robot can go. There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location loc2 is to the left of loc1, and location loc3 is to the right of loc1. Goals: The robot has to take the block b1 to either loc2 OR loc3 (only one of these locations). The human observer can not see the robot acting in the environment. Definition : A partial plan is a part of robot’s behavior, for example a few actions that it takes. Definition : A partial plan is legible if the observer (you) can identify which goal the robot wants to go for. A partial plan A is more legible than another partial plan B if the number of possible goal locations for A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step left. Question 1 : Imagine you are the human observer. Would you find such a partial plan legible? Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is legible? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.
@@
Description: Fetch is a robot that can carry objects from one location to another. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. Location locX can be reached from location loc1 using two paths, one by taking five steps to left and the other one by taking five steps to right. Goal: The robot has to pick the block b1 and take it to location locX. The human observer can not see the robot acting in the environment. Definition: A partial plan A is predictable if the observer (you) can identify if there is one possible completion (which may or may not lead to the goal). A partial plan A is more predictable than a partial plan B if the number of possible completions of A is less than B. Plan : In the robot’s partial plan, it picks block b1 and takes one step towards left. Question 1 : Imagine you are the human observer. Would you find such a partial plan predictable? Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is predictable? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.
@@
Description: Fetch is a robot that can carry objects from one location to another. Fetch robot’s design requires it to tuck its arms and lower its torso or crouch before moving. There are three locations: loc1, loc2, and loc3. Initial state: There is a block b1 at location loc1, and the robot is at location loc1 and has its hand empty. The human observer can not see the robot acting in the environment. Definition: Suppose you think the robot is trying to achieve one out of a set of of potential goals. If the agent’s behavior does not reduce the size of this set, then it is obfuscatory. For example, if you think robot is trying to achieve one of A, B, C. If it shows a behavior (partial plan) but you think it is still trying to achieve any one of A, B, C, then it is obfuscatory. Plan : Suppose the agent picks up the block b1. At loc1 is the agent is at a distance of 10 steps from loc2 and loc3. It takes 3 steps forward and is 7 steps away from loc2 as well as loc3. Question 1 : Imagine you are the human observer. Would you find such a partial plan obfuscatory? Give your answer as ‘Yes’, ‘No’ or ‘Can’t Say’ only. ——————- Question 2 : Why do you think the robot’s plan is obfuscatory? Give your response only as ’1’, ’2’ or ’3’ based on the following options and do not include any other text in your response: 1) Because the human can not see the robot acting. 2) Because the human does not like the robot’s plan. 3) Because the human can not understand the environment.